# Learning What Matters: Prioritized Concept Learning via Relative Error-driven Sample Selection

### v1

@@ Abstract
Instruction tuning has been central to the success of recent vision-language models (VLMs), but it remains expensive—requiring large scale datasets, high-quality annotations and large-compute budget. We propose PRioritized cOncept learninG via Relative Error-driven Sample Selection – PROGRESS – a data- and compute-efficient framework that enables VLMs to dynamically select what to learn next based on their evolving needs during training. At each stage, the model tracks its learning progress across skills and selects the most informative samples: those it has not already mastered and are not too difficult to learn at the current state of training. This strategy effectively controls skill acquisition and the order in which skills are learned. Specifically, we sample from skills showing the highest learning progress, prioritizing those with the most rapid improvement. Unlike prior methods, PROGRESS requires no upfront answer annotations, querying answers only on a need basis, avoids reliance on additional supervision from auxiliary VLM, or compute-heavy gradient computations for data selection. Experiments across multiple instruction-tuning datasets of varying scales demonstrate that PROGRESS consistently outperforms state-of-the-art baselines with much less data and supervision. Additionally, we show strong cross-architecture generalization to different VLMs and transferability to larger models, validating PROGRESS as a scalable solution for efficient learning. [^1] Code will be released on publication.
@@ 1 Introduction
Multimodal vision-language models (VLMs) such as GPT-4V (OpenAI et al., 2024), Gemini (Team et al., 2023), LLaVA (Liu et al., 2023b, a), and InternVL (Chen et al., 2024b) demonstrate impressive general-purpose capabilities across tasks like image comprehension and visual question answering. Much of this success stems from large-scale fine-tuning on high-quality image-text corpora, particularly visual instruction-tuning (IT) datasets (Zhang et al., 2023; Xu et al., 2024), which significantly enhance instruction-following and reasoning abilities. A growing trend in building stronger VLMs has been to simply scale up: collecting larger more diverse IT datasets with better annotations and using them to instruction-tune increasingly powerful models (Chen et al., 2023; Liu et al., 2023a).
@@
However, such pipelines are increasingly resource-intensive—annotation-heavy when relying on human-labeled supervision (e.g., bounding boxes, object tags) and monetarily costly when generating instructions via proprietary models like GPT-4 (Liu et al., 2023b, a), alongside significant computational overhead. These factors make such pipelines increasingly inaccessible to individual researchers and smaller academic labs. More importantly, it is unclear whether the entirety of these large corpora is necessary for strong VLM performance. We posit that many samples are redundant or uninformative, and that comparable results could be achieved using fewer, informative samples.
@@
To this end, we investigate how to select the most informative visual instruction-tuning (IT) samples based on the model’s own evolving learning state. We ask: Can VLMs indicate what they can most effectively learn at a give stage of training? Inspired by curriculum learning, we develop a framework in which the model periodically self-evaluates its current knowledge and identifies the skills it is ready to acquire next—those that would most benefit its learning progress. Specifically, we track the relative change in skill performance across iterations to estimate where learning improves fastest, encouraging the model to prioritize these skills. We hypothesize that this enables the VLM to actively select training samples that are most informative: those that are not already mastered by the model, and are not too difficult for the model to learn at its current stage. Overall, PROGRESS is designed to adapt to the model’s evolving learning state by helping it acquire essential skills, while also promoting diversity across selected concepts—a property crucial for capturing important modes in the data and supporting generalization.
@@
Experimental results across multiple instruction-tuning datasets of varying scale demonstrate that PROGRESS achieves up to 99–100% of the full-data performance while using only 16–20% of the labeled training data. In addition to these gains, PROGRESS offers several practical advantages over existing approaches. First, unlike static scoring-based methods(Paul et al., 2021; Coleman et al., 2019; Marion et al., 2023; Hessel et al., 2022) or concept-driven strategies that rely on additional reference VLMs (Lee et al., 2024), our method uses dynamic feedback from the model’s own learning progress to guide training. Second, while many prior methods assume full access to ground-truth annotations upfront (Lee et al., 2024; Wu et al., 2025), we operate in a more realistic setting where the training pool is initially unlabeled and answers are queried only on a need basis—drastically reducing annotation cost. Third, instead of merely selecting which samples to train on (Lee et al., 2024; Wu et al., 2025), our approach also decides when to introduce each skill—enabling curriculum-style control over both skill acquisition and learning order. Our contributions are as follows:
  • We propose PROGRESS, a dynamic, progress-driven framework for selecting the most informative samples during VLM instruction tuning—based on relative improvement across automatically discovered skills.
  • Our method achieves near full-data performance using only 16–20% supervision across multiple instruction-tuning datasets of varying scale and across different VLMs—including the widely used LLaVA-v1.5-7B. It generalizes well across architectures, showing strong results on larger-capacity models like LLaVA-v1.5-13B and newer designs such as Qwen2-VL, while consistently outperforming competitive baselines and prior data-efficient methods.
@@
  • We analyze what skills the model prioritizes and when, revealing an interesting curriculum over skill types and difficulty—offering new insights into efficient VLM training.
@@ 2 Related Work
Data Efficient Learning for VLMs. Previous approaches for efficient VLM training (see Fig.  2) typically select a coreset—i.e., a representative subset of the data—using static metrics like Clip-Score (Hessel et al., 2022) or score functions like EL2N (Paul et al., 2021), perplexity (Marion et al., 2023), entropy (Coleman et al., 2019), or train auxiliary scoring networks (Chen et al., 2024a). However, these methods perform one-time selection before training and fail to adapt to the model’s evolving needs. Score metrics often overlook important data modes, leading to poor diversity and reduced generalization (Lee et al., 2024; Maharana et al., 2025)—sometimes even underperforming random sampling (Lee et al., 2024). Other prominent work selects skill-diverse samples using reference VLMs—auxiliary models that themselves require large-scale instruction-tuning data to be effective. A notable example is COINCIDE (Lee et al., 2024), which clusters internal activations from a separately trained VLM model (e.g., TinyLLaVA (Zhou et al., 2024)) for selection. This setup requires a fully trained additional VLM, ground-truth answers for full dataset, and manual human inspection to select appropriate activations for clustering—making it resource-intensive and hard to scale. Gradient-based methods (Wu et al., 2025; Liu et al., 2024b), while principled, are computationally prohibitive—requiring substantial memory to store high-dimensional gradients and large-scale compute (often upwards of 100 GPU hours, e.g., ICONS [^2] Unpublished Concurrent work, we still reproduce and attempt to compare with it; see details in Appendix A.4. (Wu et al., 2025))—directly refuting the goal of efficient training. Some methods also assume access to explicit knowledge of the target task or its distribution in the form of labeled samples—as in ICONS (Wu et al., 2025)—which is rarely available beforehand in practical settings for general-purpose VLM training.
Figure 1: Overall Pipeline. Our framework consists of two stages: (1) Multimodal Concept Categorization, which partitions the unlabeled pool $\mathbb{U}$ into distinct skills by assigning each sample $(I,Q)$ to a specific skill cluster, and (2) Prioritized Concept Learning, where the model actively selects the most informative samples—those showing the highest improvement in its objective (e.g., accuracy or loss) relative to its prior state. Only these selected samples are annotated on a need basis, forming the labeled set $(I,Q,A)$ and used for training.
Curriculum Learning and Self-Paced Learning. Curriculum learning improves generalization by presenting data from easy to hard (Bengio et al., 2009). Self-paced extensions adapt this ordering based on model own progress (Kumar et al., 2010; Sachan & Xing, 2016). These ideas have been explored in NLP (Sachan & Xing, 2016; Mindermann et al., 2022) as well as controlled synthetic multimodal settings (Misra et al., 2017) (e.g., CLEVR) on small-scale models, using external heuristics or limiting selection to question-level. In contrast, PROGRESS scales this principle to real-world instruction tuning of large VLMs by selecting informative image-text pairs at the skill level using unsupervised representations and dynamic learning progress signals.
@@
Figure 2: Comparison with Prior Efficient Learning Methods for VLMs. Green denote desirable properties for efficient learning, while Red indicate limitations. PROGRESS satisfies all key desirable criteria while requiring only 20% data. See Appendix A for details of prior methods.
In summary, our method brings together the best of all worlds (see Fig. 2): it selects and annotates samples dynamically adapting to the model’s learning state (row 1), effectively controlling skills and order of aquisition (row 2), require no additional reference VLMs for access to internal activations, no human-in-the-loop decisions (row 3), eliminates need for explicit knowledge of the target task or its distribution (row 4), or need for compute-heavy gradient computation (row 5), promotes diverse coverage of skills (row 6), —making it a practical solution for efficient and scalable VLM training. We apply supervision strictly on a need basis to only 20% of the dataset (row 7).
@@ 3 Problem Setting and Overall Framework
Problem Setting. We now formally introduce our data-efficient learning setting for training VLMs. We denote an image by $I$, a question by $Q$, forming an image-question pair $(I,Q)\in\mathbb{U}$, where $\mathbb{U}$ is an unlabeled pool of such pairs. Unlike previous efficient learning methods, we do not assume access to the corresponding answers $A\in\mathbb{A}$ for all pairs in $\mathbb{U}$, and thus refer to this pool as unlabeled. The learner is provided with: (1) the unlabeled pool $\mathbb{U}$; and (2) a fixed answer budget $b$, specifying the maximum number of pairs from $\mathbb{U}$ for which it can query an answer $A\in\mathbb{A}$ and use for training, where $|\mathbb{A}|=b\ll|\mathbb{U}|$. The goal is to learn a vision-language model $\mathtt{VLM}(A\mid I,Q)$ that can accurately predict an answer for a new image-question pair, while only using $b$ selected and labeled samples during training. The central challenge lies in identifying the most informative $(I,Q)$ pairs to annotate within the constrained budget $b$, such that the resulting model trained on these $(I,Q,A)$ pairs performs comparably to one trained on the fully labeled dataset.
Overall Framework. Our overall framework for efficient training of VLMs is shown in Figure 1. We employ a two-stage pipeline:
@@
    Multimodal Concept Categorization. Given an unlabeled data pool $\mathbb{U}$ containing image-question pairs $(I,Q)\in\mathbb{U}$, we first partition $\mathbb{U}$ into a distinct set of $K$ skills, assigning each sample $(I,Q)$ to a specific skill. This categorization enables tracking the model’s progress on individual skills and supports a self-paced training strategy where the model’s own learning signals determine which skills to prioritize next.
@@
Overall, our model dynamically selects diverse, informative samples throughout training, in alignment with its evolving learning needs. To ensure that skill-level performance estimates are credible at the start of training—when the model is still untrained—we begin with a brief warmup phase (see details in Appendix A). This warmup, along with subsequent samples selected through our prioritized strategy, together make up the total annotation budget of $b$, ensuring that supervision in the form of $(I,Q,A)$ never exceeds the specified budget. We validate that using the warmup set alone to train the model yields significantly worse performance compared to our proposed strategy—highlighting the importance of progress-driven sampling. Our framework allows the model to be trained with much less data and supervision, while controlling both skill acquisition and learning order.
@@ 3.1 Multimodal Concept Categorization
We begin by identifying diverse skills from the unlabeled data pool through a fully unsupervised concept categorization module that partitions $\mathbb{U}$ into $K$ skill clusters using spherical k-means.
Figure 3: Cluster Visualization. Clustering with multimodal DINO-BERT features ensures purer skill clusters with higher intra-cluster and lower inter-cluster similarity compared to uni-modal partitioning
Each sample $(I,Q)\in\mathbb{U}$ is assigned to a cluster based on cosine similarity from multimodal concatenated self-supervised DINO (Oquab et al., 2024) (for image $I$) and BERT (Devlin et al., 2019) (for text question $Q$) features. Jointly leveraging both modalities yields purer clusters with higher intra-cluster and lower inter-cluster similarity compared to unimodal partitioning (see Fig. 3)—enabling accurate tracking of skill-level progress during training. Unlike COINCIDE (closest best performing prior work) (Lee et al., 2024), our concept categorization framework is simpler, unsupervised, and more scalable. COINCIDE requires activations from fully trained additional VLM, ground-truth answers for full dataset, and human inspection to select appropriate activations for skill clustering. In contrast, our categorization is fully automated and practical, requires no labels, or manual introspection.
@@
Our goal is to guide the VLM to prioritize skills it can most readily learn and improve upon. Since human intuition about task difficulty may not align with model’s difficulty in its feature and hypothesis space (Sachan & Xing, 2016), we adopt a self-paced strategy where the model’s own learning progress determines what to learn next. Inspired by curriculum learning (Kumar et al., 2010; Sachan & Xing, 2016), we select the most informative samples—those that yield the greatest improvement in the model’s objective (e.g., accuracy or loss) relative to its prior state.
@@
Formally, given an unlabeled pool $\mathbb{U}=\{(I,Q)\}$ partitioned into skill clusters $\mathcal{C}=\{C_{1},C_{2},\dots,C_{K}\}$, we define the model’s learning state at step $t$ by its accuracy $\text{Acc}_{k}^{(t)}$ on each cluster $k$, computed over the training data seen by the model so far. The relative change in performance across steps quantifies learning progress per skill, which is then used to guide sample selection. We compute the expected accuracy improvement for each skill cluster between step $t$ and $t-\gamma$:
$\Delta_{k}=\frac{\text{Acc}_{k}^{(t)}-\text{Acc}_{k}^{(t-\gamma)}}{\text{Acc}_{k}^{(t-\gamma)}+\epsilon}$ (1)
@@
where $\epsilon$ ensures numerical stability. The score $\Delta_{k}$ captures how rapidly the model is improving on skill cluster $k$, serving as a proxy for sample informativeness. By prioritizing high $\Delta_{k}$ clusters , the model focuses on skills it can improve on most rapidly—thereby enforcing a self-paced curriculum that dynamically adapts to the model’s learning state (Sachan & Xing, 2016)—controlling both the acquisition of skills and the order in which they are learned. In addition to the selection strategy in Eqn 1, we sample $\delta\%$ of points at random to encourage exploration of new or underrepresented skills in dataset, following prior curriculum learning work (Kumar et al., 2010; Misra et al., 2017).
@@
Only selected samples are annotated, forming the labeled set $(I,Q,A)$ for training. This need-based annotation strategy avoids the costly requirement of full supervision used in prior coreset methods (such as COINCIDE (Lee et al., 2024)), offering a more scalable and efficient training paradigm.
@@
However, naively selecting samples from only the highest-improvement cluster can hurt diversity by concentrating on a narrow skill set and leading to mode collapse—an issue known to degrade performance in prior work (Lee et al., 2024). To mitigate this, we propose to sample from multiple high $\Delta_{k}$ clusters in proportion to their relative improvement using a temperature-controlled softmax:
$p_{k}=\frac{\exp(\Delta_{k}/\tau)}{\sum_{j=1}^{K}\exp(\Delta_{j}/\tau)}$ (2)
@@
Here, $p_{k}$ is the probability of sampling from cluster $k$, and $\tau$ controls the sharpness of the distribution. Lower $\tau$ emphasizes top clusters but risks mode collapse by repeatedly sampling from a narrow skill set (higher informativeness, lower diversity); higher $\tau$ promotes broader sampling and better skill coverage. This balance between informativeness and diversity is critical for effective and robust learning (see ablation Fig. 4). The sampling budget at given step $t$ is then allocated proportionally to $p_{k}$, and only the selected samples are annotated as $(I,Q,A)$ triplets for training.

### v2

@@ Abstract
Instruction tuning has been central to the success of recent vision-language models (VLMs), but it remains expensive—requiring large scale datasets, high-quality annotations and large-compute budget. We propose PRioritized cOncept learninG via Relative Error-driven Sample Selection – PROGRESS – a data- and compute-efficient framework that enables VLMs to dynamically select what to learn next based on their evolving needs during training. At each stage, the model tracks its learning progress across skills and selects the most informative samples: those it has not already mastered and are not too difficult to learn at the current state of training. This strategy effectively controls skill acquisition and the order in which skills are learned. Specifically, we sample from skills showing the highest learning progress, prioritizing those with the most rapid improvement. Unlike prior methods, PROGRESS requires no upfront answer annotations, querying answers only on a need basis, avoids reliance on additional supervision from auxiliary VLM, or compute-heavy gradient computations for data selection. Experiments across multiple instruction-tuning datasets of varying scales demonstrate that PROGRESS consistently outperforms state-of-the-art baselines with much less data and supervision. Additionally, we show strong cross-architecture generalization to different VLMs and transferability to larger models, validating PROGRESS as a scalable solution for efficient learning. [^1] Project website: https://mylittlechange.github.io/PROGRESS_web/
@@ 1 Introduction
Multimodal vision-language models (VLMs) such as GPT-4V [35], Gemini [41], LLaVA [25, 24], and InternVL [7] demonstrate impressive general-purpose capabilities across tasks like image comprehension and visual question answering. Much of this success stems from large-scale fine-tuning on high-quality image-text corpora, particularly visual instruction-tuning (IT) datasets [48, 47], which significantly enhance instruction-following and reasoning abilities. A growing trend in building stronger VLMs has been to simply scale up: collecting larger more diverse IT datasets with better annotations and using them to instruction-tune increasingly powerful models [5, 24].
@@
However, such pipelines are increasingly resource-intensive—annotation-heavy when relying on human-labeled supervision (e.g., bounding boxes, object tags) and monetarily costly when generating instructions via proprietary models like GPT-4 [25, 24], alongside significant computational overhead. These factors make such pipelines increasingly inaccessible to individual researchers and smaller academic labs. More importantly, it is even unclear whether the entirety of these large corpora is necessary for strong VLM performance. We posit that many samples are redundant or uninformative, and that comparable results could be achieved using fewer, informative samples.
@@
To this end, we investigate how to select the most informative visual instruction-tuning (IT) samples based on the model’s own evolving learning state. We ask: Can VLMs indicate what they can most effectively learn at a give stage of training? Inspired by curriculum learning, we develop a framework in which the model periodically self-evaluates its current knowledge and identifies the skills it is ready to acquire next—those that would most benefit its learning progress. Specifically, we track the relative change in skill performance across iterations to estimate where learning improves fastest, encouraging the model to prioritize these skills. We hypothesize that this enables the VLM to actively select training samples that are most informative: those that are not already mastered by the model, and are not too difficult, allocating its limited budget to those offering the highest learning benefit and thereby improving overall training efficiency. Overall, PROGRESS adapts to the model’s evolving learning state and helps acquire essential skills while balancing informativeness of selected samples with diversity of concepts—a property crucial for covering important data modes and supporting generalization.
@@
Experimental results across multiple instruction-tuning datasets of varying scale demonstrate that PROGRESS achieves up to 99–100% of full-data performance using only 16–20% of the labeled training data, while also being faster in overall training time—including all self-evaluation overhead. In addition to these gains, PROGRESS offers several practical advantages over existing approaches. First, unlike static scoring-based methods [36, 9, 31, 13] or concept-driven strategies that rely on additional reference VLMs [19], our method uses dynamic feedback from the model’s own learning progress to guide training. Second, while many prior methods assume full access to ground-truth annotations upfront [19, 45], we operate in a more realistic setting where the training pool is initially unlabeled and answers are queried only on a need basis—drastically reducing annotation cost. Third, instead of merely selecting which samples to train on [19, 45], our approach also decides when to introduce each skill—enabling curriculum-style control over both skill acquisition and learning order. Our contributions are as follows:
  • We propose PROGRESS, a dynamic, progress-driven framework that selects the most informative samples for VLM instruction tuning by leveraging the model’s own learning progress across automatically discovered multimodal concepts—without relying on auxiliary VLMs, manual heuristics, or full-data supervision.
  • PROGRESS is data and label efficient – achieves near full-data performance (99–100%) using only 16–20% labeled data, generalizing across datasets, architectures, and model scales (LLaVA-v1.5-7B/13B, Qwen2-VL).
  • PROGRESS is also faster in total training time, including the self-evaluation overhead.
@@
  • We analyze what the model learns and when, revealing an interesting curriculum over concept types and difficulty, offering new insights into efficient VLM training.
@@ 2 Related Work
Data Efficient Learning for VLMs. Prior approaches to efficient VLM training (see Fig. 1) typically select a coreset—i.e., a representative subset of the data for training—using static metrics such as CLIP-Score [13], EL2N [36], perplexity [31], entropy [9], or by training auxiliary scoring networks [6]. However, these methods perform one-time sample selection before training and cannot adapt to the model’s evolving needs. Static score metrics also miss important data modes, leading to poor diversity and reduced generalization [19, 30], and in some cases even underperform random selection [19]. Gradient-based methods [45, 27], while more principled, are computationally prohibitive—requiring large memory to store high-dimensional gradients and hundreds of GPU hours (e.g., ICONS [^2] Unpublished concurrent work; we reproduce and compare in Appendix 6.4 [45])—contradicting the very goal of efficient training. Some also assume access to explicit knowledge of the target task or distribution through labeled samples, as in ICONS [45], which is rarely realistic in general-purpose VLM settings. More recently, prominent work has explored selection using additional reference VLMs—auxiliary models that themselves require large-scale instruction tuning. A notable example is COINCIDE [19], which extracts internal activations from a separately trained VLM (e.g., TinyLLaVA [50]) to guide coreset selection. However, COINCIDE exhibits several limitations: it requires a fully trained additional auxiliary VLM, performs static one-time selection without controlling the order of skills being learned, needs ground-truth annotations for the entire dataset, and requires manual human inspection to first select appropriate activations—all of which make it resource-intensive and difficult to scale.
@@
Figure 1: Comparison with Prior Efficient Learning Methods for VLMs. Green denote desirable properties for efficient learning, while Red indicate limitations. PROGRESS satisfies all key desirable criteria while requiring only 20% data. See Appendix 6.1 for details of prior approaches.
Figure 2: Overall Pipeline. Our framework consists of two stages: (1) Multimodal Concept Categorization, which partitions the unlabeled pool $\mathbb{U}$ into distinct skills by assigning each sample $(I,Q)$ to a specific skill cluster, and (2) Prioritized Concept Learning, where the model actively selects the most informative samples—those showing the highest improvement in its objective (e.g., accuracy or loss) relative to its prior state. We query annotations for only these selected samples on a need basis, forming labeled set $(I,Q,A)$, which is used for training.
Curriculum Learning, Self-Paced Learning, and Active Learning. Curriculum learning improves generalization by ordering data from easy to hard [3], while self-paced learning adapts this order based on the learner’s progress [18, 38]. These ideas have been explored in NLP [38, 33] and in controlled synthetic multimodal settings [34], but typically with small-scale models and externally designed heuristics. To the best of our knowledge, PROGRESS is the first method to show how these principles can be used to train large-scale VLMs efficiently. PROGRESS shares the overall goal of active learning in selecting informative samples, but rather than relying on instance-level uncertainty, it prioritizes samples where the model shows the greatest relative improvement—yielding a progress-driven, curriculum-like ordering instead.
Key Differences Over Prior Methods. PROGRESS introduces a novel framework that uses the model’s own learning progress to select the most informative samples for data-efficient training of large VLMs, in contrast to COINCIDE, which relies on a separately trained auxiliary VLM. PROGRESS dynamically adapts to the model’s evolving learning state, explicitly controlling both which skills are acquired and the order in which they are learned (rows 1–2 in Fig. 1), whereas COINCIDE performs a static, one-shot selection that ignores acquisition order. In PROGRESS supervision is applied strictly on a need-to-label basis, requiring annotations for only $\sim$20% of the dataset (row 7), while COINCIDE assumes full-data labels. Moreover, PROGRESS requires no target-task knowledge (row 4), avoids compute-heavy gradient estimation (row 5), and maintains diverse skill coverage (row 6). To the best of our knowledge, PROGRESS is the first method to use a model’s own signal to drive a self-paced, progress-driven curriculum for efficient large-VLM training—focusing learning where progress is greatest while jointly controlling both skill acquisition and learning order.
@@ 3 Problem Setting and Overall Framework
Problem Setting. We now formally introduce the data-efficient learning setting for training VLMs. We denote an image by $I$, a question by $Q$, forming an image-question pair $(I,Q)\in\mathbb{U}$, where $\mathbb{U}$ is an unlabeled pool of such pairs. Unlike previous efficient learning methods, we do not assume access to the corresponding answers $A\in\mathbb{A}$ for all pairs in $\mathbb{U}$, and thus refer to this pool as unlabeled. The learner is provided with: (1) the unlabeled pool $\mathbb{U}$; and (2) a fixed answer budget $b$, specifying the maximum number of pairs from $\mathbb{U}$ for which it can query an answer $A\in\mathbb{A}$ and use for training, where $|\mathbb{A}|=b\ll|\mathbb{U}|$. The goal is to learn a vision-language model $\mathtt{VLM}(A\mid I,Q)$ that can accurately predict an answer for a new image-question pair, while only using $b$ selected and labeled samples during training. The central challenge lies in identifying the most informative $(I,Q)$ pairs to annotate within the constrained budget $b$, such that the resulting model trained on these $(I,Q,A)$ pairs performs comparably to one trained on the fully labeled dataset. This setup follows standard setting used in prior data-efficient learning [19, 45, 6]
Overall Framework. Our overall framework is shown in Figure 2. We employ a two-stage pipeline:
@@
    Multimodal Concept Categorization. Given an unlabeled data pool $\mathbb{U}$ with image-question pairs $(I,Q)\in\mathbb{U}$, we first partition $\mathbb{U}$ into $K$ concept clusters in a fully unsupervised manner, assigning each sample $(I,Q)$ to a specific concept. This enables tracking the model’s progress on individual concepts and supports a self-paced training strategy where the model’s own learning signals determines which concepts to prioritize next.
@@
To obtain reliable skill-level performance estimates at the start of training—when the model is still untrained—we include a brief warmup phase, consistent with prior work. We adopt simple sampler, selecting samples from our unsupervised concept clusters with probability: $P_{i}\propto\exp\left({S_{i}}/{\tau D_{i}}\right)$ where $S_{i}$ denotes the cluster’s transferability and $D_{i}$ its density [19] (see details in Appendix 6.3). The warmup samples, together with those selected by our prioritized strategy, constitute the total annotation budget $b$, ensuring that the labeled set $(I,Q,A)$ never exceeds this budget. Training with the warmup set alone yields substantially worse performance, highlighting the importance of our core contribution – progress-driven sampling (Table 4).
@@ 3.1 Multimodal Concept Categorization
We first get a set of concept clusters from the unlabeled instruction-tuning pool $U$. Instead of manually defining categories, we aim to obtain coherent visual–linguistic groupings where samples share similar image content and question type; we refer to these groupings as concepts or skills. For each image–question pair $(I,Q)$, we extract self-supervised features using a frozen DINO vision encoder $f_{\text{img}}(I)$ and a BERT question encoder $f_{\text{text}}(Q)$, concatenate and L2-normalize them: $z(I,Q)=\operatorname{norm}\big([f_{\text{img}}(I);\;f_{\text{text}}(Q)]\big).$ We then apply spherical k-means to partition the pool into $K$ concept clusters $\{C_{1},\dots,C_{K}\}$, each serving as a concept group for tracking learning progress. Using both modalities yields more coherent and interpretable clusters than unimodal variants; as shown in Fig.3 and Appendix 8.1 (Word Cloud Visualization), these groups often align with abilities such as object grounding, OCR, programming and multilingual tasks. Unlike COINCIDE—our closest prior competitor—which requires activations from an additional VLM, full annotations, and manual inspection, our categorization is fully unsupervised & model-agnostic, requiring no annotations, auxiliary models, or human intervention.
Figure 3: Cluster Visualization. Clustering with multimodal DINO-BERT features ensures purer skill clusters with higher intra-cluster and lower inter-cluster similarity compared to uni-modal partitioning. See Word Cloud Visualization Appendix 8.1.
@@
Our goal is to guide the VLM toward the skills it can most effectively learn at its current stage of training. To use the annotation budget efficiently, we aim to select samples that offer high learning benefit—avoiding those that are too easy (redundant, low gain) or too hard (beyond the model’s current capacity). Concretely, we leverage the model’s own learning progress to identify informative samples: those that yield the largest improvement in the model’s objective (e.g., accuracy or loss) relative to its prior state.
@@
Formally, given an unlabeled pool $\mathbb{U}=\{(I,Q)\}$ partitioned into concept clusters $\mathcal{C}=\{C_{1},\dots,C_{K}\}$, we track the model’s learning state at step $t$ using its accuracy $\text{Acc}_{k}^{(t)}$ on each cluster $k$, computed over the labeled training data observed so far. The relative change in accuracy quantifies learning progress and drives sample selection. For each cluster, we compute improvement between steps $t$ and $t-\gamma$:
$\Delta_{k}=\frac{\text{Acc}_{k}^{(t)}-\text{Acc}_{k}^{(t-\gamma)}}{\text{Acc}_{k}^{(t-\gamma)}+\epsilon},$ (1)
@@
where $\epsilon$ ensures stability. The score $\Delta_{k}$ measures how rapidly the model is improving on cluster $k$ and serves as a proxy for sample informativeness. Prioritizing high-$\Delta_{k}$ clusters yields a self-paced curriculum that adapts to the model’s evolving capabilities, focusing learning where progress is greatest - controlling both the acquisition of skills and the order in which they are learned. For exploration, we additionally sample $\delta\%$ of points uniformly at random, following prior curriculum-learning work [18, 34].
@@
Annotations are queried only for the selected samples, forming the labeled set $(I,Q,A)$. This need-based annotation strategy removes the requirement for full-dataset supervision used in prior coreset methods [19], enabling far more scalable and efficient training.
@@
However, selecting samples only from the highest-improvement cluster hurt diversity by concentrating on a narrow skill set and leads to mode collapse—an issue also noted in prior work [19]. To mitigate this, we propose to sample from multiple high-$\Delta_{k}$ clusters using a temperature-controlled softmax:
$p_{k}=\frac{\exp(\Delta_{k}/\tau)}{\sum_{j=1}^{K}\exp(\Delta_{j}/\tau)}.$ (2)
@@
Here, $p_{k}$ is the probability of sampling cluster $k$, and $\tau$ controls the sharpness of the distribution. Lower $\tau$ concentrates sampling on top-improving clusters (high informativeness but low diversity), while higher $\tau$ broadens sampling & improves skill coverage. Balancing informativeness vs. diversity is essential for learning ( Fig. 5). At each step $t$, the sampling budget is allocated proportionally to $p_{k}$, & only selected samples are annotated as $(I,Q,A)$ for training.
Table 1: Comparison of coreset selection techniques for training LLaVA-v1.5 on the LLaVA-665K dataset using 20% sampling ratio. AL Only refers to methods which align with active learning principles. AL & CL refers to methods which use combination of active and curriculum learning principles. Methods highlighted in orange require additional reference VLMs and 100% dataset annotations for coreset selection, while methods highlighted in light green do not require either and thus considered better. The benchmark results are highlighted with best and second best within the respective categories (i.e, with and without utilizing additional information). The best and the second best relative score are in bold and underlined, respectively. Reported numbers are averaged over 3 runs.
[…]
22footnotetext: Reproduced with official code.